OpenStudy (openstudyrocks5*):

What is the solution to (4 x 10^3) x (3 x 10^4) written in scientific notation? EXPLAIN

OpenStudy (openstudyrocks5*):

No not really.......

OpenStudy (openstudyrocks5*):

@skullpatrol What are you doing? Why do you keep putting stuff then deleting?

OpenStudy (jakob):

First off I am learning with you so excuse me if I am annoying. What is scientific notation?

OpenStudy (openstudyrocks5*):

Um....

OpenStudy (openstudyrocks5*):

that what you meant?

OpenStudy (jakob):

So first off it kinda looks like we are going to have to use PEMDAS..... Do you know what PEMDAS is?

OpenStudy (openstudyrocks5*):

no, and im getting more confused than i was before XD

OpenStudy (jakob):

Hold on this is a moderator

OpenStudy (openstudyrocks5*):

What? xD

OpenStudy (jakob):

she can be more of service than I can

OpenStudy (jakob):

or he

OpenStudy (openstudyrocks5*):

or "The cat"

OpenStudy (jakob):

Lol

OpenStudy (welshfella):

Scientific notation is of the form N * 10^n Wher N is a number from 1 to 10 ( not including 10) and n is an integer ( except 0)

OpenStudy (welshfella):

step 1 multiply 3 by 4 step 2 multiply 10^3 by 10^4 (using laws of exponents)

OpenStudy (welshfella):

can you do that?

OpenStudy (openstudyrocks5*):

i don't think i can do step 2

OpenStudy (welshfella):

10^3 * 10^4 = 10^(3+4) = 10^7 when you multiply you ADD the exponents.

OpenStudy (openstudyrocks5*):

Oh, That makes sense. ok.....

OpenStudy (welshfella):

so we have 12 * 10^7 but athat's not correct scientific notation beacuse 12 is not between 1 and 10

OpenStudy (openstudyrocks5*):

Yeah..

OpenStudy (welshfella):

do convert to the correct form we divide the first part by 10 and multiply the exponent part by 10

OpenStudy (openstudyrocks5*):

by first part you mean 12 right?

OpenStudy (welshfella):